OpenStudy (anonymous):

square root of (4t^2 +1)

OpenStudy (unklerhaukus):

\[\sqrt{4t^2+1}\] \[\sqrt{t^2(4+1/t^2)}\] \[=t\sqrt{4+\frac{1}{t^2}}\]

OpenStudy (anonymous):

t=i/2 or t=-i/2